OpenStudy (anonymous):

Calculate the discriminant. x^2 + 2x – 2 = 0

OpenStudy (anonymous):

I got -4 , and I was wrong :(

OpenStudy (whpalmer4):

What are you using for the formula for the discriminant, and what are you using for your values of \(a,b,c\)?

OpenStudy (anonymous):

b^2 -4ac b=2, a=1, c=2

OpenStudy (whpalmer4):

well, one of those is not correct. Compare your equation: \[x^2+2x-2 = 0\]\[ax^2+bx+c=0\]

OpenStudy (anonymous):

The -2 and +C ?

OpenStudy (whpalmer4):

right. c = -2, not c = 2

OpenStudy (anonymous):

So that's were I got it wrong. Thank you :)

OpenStudy (whpalmer4):

what's your new, improved answer?

OpenStudy (anonymous):

12

OpenStudy (whpalmer4):

very good!